# On the ring of cooperations for Real Hermitian K-theory

### v1

@@ 4.2. The algebraic Atiyah–Hirzebruch spectral sequence
$0\to\Sigma^{2(k-1),k-1}B_{0}^{F}(\tfrac{k-1}{2})\otimes_{\mathbb{M}_{2}^{F}}B_{0}^{F}(1)\to B_{0}^{F}(k)\to B_{1}^{F}(\tfrac{k-1}{2}-1)\otimes_{\mathbb{M}_{2}^{F}}(\mathcal{A}(1)\operatorname{/\!/}\mathcal{A}(0))_{F}^{\vee}\to 0$
As before, the long exact sequence induced by $\text{Ext}_{\mathcal{A}(1)^{\vee}_{F}}(-)$ splits into short exact sequences. The cokernels are again a simple computation, but the kernels require a more delicate analysis. Indeed, to compute $\text{Ext}_{\mathcal{A}(1)^{\vee}_{F}}\left(B_{0}^{F}(\tfrac{k-1}{2})\otimes_{\mathbb{M}_{2}^{F}}B_{0}^{F}(1)\right)$, we use another aAHSS. This spectral sequence comes from applying $\text{Ext}_{\mathcal{A}(1)_{F}^{\vee}}\left(B_{0}^{F}(\tfrac{k-1}{2})\otimes_{\mathbb{M}_{2}^{F}}-\right)$ to the previous cellular filtration on $B_{0}^{F}(1)$. The $E_{1}$-page is given by
$E_{1}=\text{Ext}_{\mathcal{A}(1)^{\vee}}(B_{0}^{F}(\tfrac{k-1}{2}))\otimes_{\mathbb{M}_{2}^{F}}\mathbb{M}_{2}^{F}\{[1],[\overline{\xi}_{1}],[\overline{\tau}_{1}]\}.$
@@
[…]
Remark 4.8.
One can also use a similar analysis to analyze the $n$-line of the kq-resolution, which takes the form
$E_{1}^{*,n,*}=\pi^{F}_{*+n,*}(\text{kq}\otimes\overline{\text{kq}}^{\otimes n}).$
We may compute the $n$-line of the kq-resolution by a motivic Adams spectral sequence. The Künneth spectral sequence for $\textup{H}_{**}(\textup{kq}\otimes\overline{\textup{kq}}^{\otimes n})$ takes the form
$E_{2}=\text{Tor}^{\mathbb{M}_{2}^{F}}(\textup{H}_{**}(\textup{kq}),\textup{H}_{**}(\overline{\textup{{kq}}}^{\otimes n}))\implies\textup{H}_{**}(\textup{kq}\otimes\overline{\textup{kq}}^{\otimes n}).$
Using Proposition 2.20, one sees that this collapses on the $E_{2}$-page. This implies that
$\text{H}_{**}(\text{kq}\otimes\overline{\text{kq}}^{\otimes n})\cong\text{H}_{**}(\text{kq})\otimes_{\mathbb{M}_{2}^{F}}\text{H}_{**}(\overline{\text{kq}})^{\otimes n}.$
After using the change of rings isomorphism on $\text{H}_{**}(\textup{kq})$, the $\textbf{mASS}(\textup{kq}\otimes\overline{\textup{kq}}^{\otimes n})$ takes the form
$E_{2}=\bigoplus_{I\in\mathcal{I}_{n}}\Sigma^{4|I|,2|I|}\text{Ext}_{\mathcal{A}(1)^{\vee}_{F}}(B_{0}^{F}(I))\implies\pi_{*+n,*}^{F}(\textup{kq}\otimes\overline{\textup{kq}}^{\otimes n}),$
where $\mathcal{I}_{n}=\{I=(i_{1},\ldots,i_{n}):i_{j}\geq 0\text{ for all }1\leq j\leq n\}$, and if $I=(i_{1},\ldots,i_{n})$, then $|I|=i_{1}+\cdots+i_{n}$ and $B_{0}^{F}(I)=B_{0}^{F}(i_{1})\otimes\cdots\otimes B_{0}^{F}(i_{n})$. Once one has determined the structure of the groups $\text{Ext}_{\mathcal{A}(1)^{\vee}_{F}}(B_{0}^{F}(k))$ through the process of computing the ring of cooperations, one can study this spectral sequence in detail to compute the $E_{1}$-page of the kq-resolution.
@@ 5. Computations in $\text{SH}(\mathbb{C})$
In this section, we review Culver–Quigley’s [CQ21] computation of $\pi_{**}^{\mathbb{C}}(\textup{kq}\otimes\textup{kq})$. This serves both as an example of the methods described in Section 2 and as a place to recall results which will be useful in the sequel via base change. There are a few places where our arguments differ from the ones given in [CQ21]; we will explicitly indicate when this is the case.
@@ 5.1. Background
Let $\mathbb{M}_{2}^{\mathbb{C}}=\mathbb{F}_{2}[\tau]$ denote the mod-2 motivic homology of a point [Voe03a, Corollary 6.10], where $|\tau|=(0,-1).$ Recall that as $\rho=0$ in $\mathbb{M}_{2}^{\mathbb{C}}$, the dual Steenrod algebra is an honest Hopf algebra (see Remark 2.9). For the reader’s convenience, we review its structure.
Theorem 5.1.
[Voe03b, Theorem 12.6] The dual motivic Steenrod algebra $\mathcal{A}^{\vee}_{\mathbb{C}}=\pi_{**}^{\mathbb{C}}(\textup{H}\mathbb{F}_{2}\otimes\textup{H}\mathbb{F}_{2})$ is the commutative Hopf algebra given by
@@
There is a well understood relationship between $\mathcal{A}_{\mathbb{C}}^{\vee}$ and the classical dual Steenrod algebra $\mathcal{A}^{\vee}$. Complex Betti realization gives a map $\text{Be}^{\mathbb{C}}:\mathcal{A}^{\vee}_{\mathbb{C}}\to\mathcal{A}^{\vee}$ determined by
@@
This extends to an isomorphism in Ext [DI10]:
$\text{Ext}_{\mathcal{A}_{\mathbb{C}}^{\vee}}(\mathbb{M}_{2}^{\mathbb{C}})\otimes_{\mathbb{M}_{2}^{\mathbb{C}}}\mathbb{M}_{2}^{\mathbb{C}}[\tau^{-1}]\cong\text{Ext}_{\mathcal{A}^{\vee}}(\mathbb{F}_{2})\otimes\mathbb{F}_{2}[\tau^{\pm 1}].$
This observation is at the heart of the theory of synthetic spectra [Pst23], and will allow us to determine motivic information from classical information.
In addition to an understanding of the homology $\textup{H}_{**}(\textup{kq})\cong(\mathcal{A}\operatorname{/\!/}\mathcal{A}(1))_{\mathbb{C}}^{\vee}$, we also have a computation of the homotopy groups $\pi_{**}^{\mathbb{C}}(\textup{kq})$ [IS11].
Theorem 5.3.
[…]
$E_{2}^{s,f,w}=\textup{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(\mathbb{M}_{2}^{\mathbb{C}})=\frac{\mathbb{M}_{2}^{\mathbb{C}}[h_{0},h_{1},a,b]}{(h_{0}h_{1},\tau h_{1}^{3},h_{1}a,a^{2}=h_{0}^{2}b)},$
where $|h_{0}|=(0,1,0)$, $|h_{1}|=(1,1,1)$, $|a|=(4,3,2)$ and $|b|=(8,4,4)$. Moreover, the spectral sequence collapses at $E_{2}$ with no extensions, hence there is an isomorphism
$\pi_{**}^{\mathbb{C}}\textup{kq}\cong\frac{\mathbb{Z}_{2}^{\wedge}[\tau,h_{1},a,b]}{(2h_{1},\tau h_{1}^{3},h_{1}a,a^{2}=4b)},$
@@
[…]
$E_{2}^{s,f,w}=\bigoplus_{k\geq 0}\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}^{s,f,w}(\Sigma^{4k,2k}B_{0}^{\mathbb{C}}(k)),$
where $B_{0}^{\mathbb{C}}(k)$ are the $\mathbb{C}$-motivic integral motivic Brown–Gitler comodules. Therefore, we must compute the trigraded groups $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(B_{0}^{\mathbb{C}}(k))$ for $k\geq 0.$ Since $B_{0}^{\mathbb{C}}(0)=\mathbb{M}_{2}^{\mathbb{C}}$, the first summand is computed by 5.3. We move then to compute $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(B_{0}^{\mathbb{C}}(1))$ by the algebraic Atiyah–Hirzebruch spectral sequence outlined in Section 4.2.
5.2. $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(B_{0}^{\mathbb{C}}(1)^{\otimes i})$
@@
$E_{1}^{s,f,w,a}=\text{Ext}^{s,f,w}_{\mathcal{A}(1)_{\mathbb{C}}^{\vee}}(\mathbb{M}_{2}^{\mathbb{C}})\otimes_{\mathbb{M}_{2}^{\mathbb{C}}}\mathbb{M}^{\mathbb{C}}_{2}\{[1],[\overline{\xi}_{1}],[\overline{\tau}_{1}]\}\implies\text{Ext}^{s,f,w}_{\mathcal{A}(1)_{\mathbb{C}}^{\vee}}(B_{0}^{\mathbb{C}}(1)),$
with differentials $d_{r}:E_{r}^{s,f,w,a}\to E_{r}^{s-1,f+1,w,a-r}$. We will let $\alpha[i]$ denote the copy of a class $\alpha\in\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(\mathbb{M}_{2}^{\mathbb{C}})$ in Atiyah–Hirzebruch filtration $i$. We have discussed in Section 4.2 the differentials for this spectral sequence. Namely, the $d_{1}$-differential is only nonzero on Atiyah–Hirzebruch filtration 3, where
@@
There is a potential nonzero third differential $d_{3}:E_{3}^{s,f,w,3}\to E_{3}^{s-1,f+1,w,0}$. However, for degree reasons we have that $\langle\alpha,h_{0},h_{1}\rangle=0$ for all remaining classes $\alpha[3]\in E_{3}^{s,f,w,3}$, hence this differential is 0. Thus the spectral sequence collapses on $E_{3}=E_{\infty}$.
We depict the $E_{1}$-page in Fig. 3, the $E_{2}$-page in Fig. 4, and the $E_{3}=E_{\infty}$-page in Fig. 5. Charts are written in $(s,f)$-grading, with motivic weight and Atiyah–Hirzebruch filtration suppressed. A black $\bullet$ represents $\mathbb{M}_{2}^{\mathbb{C}}$. A hollow $\circ$ represents $\mathbb{F}_{2}$. A black vertical line represents multiplication by $h_{0}$. A diagonal line represents multiplication by $h_{1}$. Differentials are blue, linear with respect to the underlying $\text{Ext}_{\mathcal{A}(1)_{\mathbb{C}}^{\vee}}(\mathbb{M}_{2}^{\mathbb{C}})$-module structure, and preserve motivic weight. A dashed blue line represents a differential where either source or target is $\tau$-torsion.
Figure 3. The $E_{1}$-page of the $\textbf{aAHSS}(B_{0}^{\mathbb{C}}(1))$
Figure 4. The $E_{2}$-page of the aAHSS($B_{0}^{\mathbb{C}}(1)$)
@@
The hidden extensions on the $E_{\infty}$-page can be resolved using complex Betti realization. Recall from Section 2.1 that this is a functor $\text{Be}^{\mathbb{C}}:\text{SH}(\mathbb{C})\to\text{Sp}$. For $X\in\text{SH}(\mathbb{C})$, this gives a map in homotopy
$\pi_{s,w}^{\mathbb{C}}X\to\pi_{s}\text{Be}^{\mathbb{C}}X.$
In particular, there is a map of dual Steenrod algebras $\mathcal{A}^{\vee}_{\mathbb{C}}\to\mathcal{A}^{\vee}$ sending $\bar{\xi}_{i}\mapsto\bar{\xi}_{i}^{2}$ and $\bar{\tau}_{i}\mapsto\bar{\xi}_{i}$. Additionally, there is an isomorphism of $\mathcal{A}(1)_{\mathbb{C}}^{\vee}$-comodules
$(B_{0}^{\mathbb{C}}(1))[\tau^{-1}]\cong B_{0}(1)[\tau^{\pm 1}],$
where $B_{0}(1)$ is the classical Brown–Gitler comodule. This induces a map of Ext groups
$\text{Ext}^{s,f,w}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(B_{0}^{\mathbb{C}}(k))\to\text{Ext}_{\mathcal{A}(1)^{\vee}}^{s,f}(B_{0}(k)),$
and allows one to lift hidden extensions from the classical case in [Mah81] to the $\mathbb{C}$-motivic case.
@@
Corollary 5.4.
[CQ21, Corollary 3.31] There is an isomorphism:
$\textup{Ext}_{\mathcal{A}(1)_{\mathbb{C}}^{\vee}}(\mathbb{M}_{2}^{\mathbb{C}},B_{0}^{\mathbb{C}}(1))\cong\pi^{\mathbb{C}}_{**}(\textup{ksp}).$
@@
[…]
$E_{2}=\text{Ext}_{\mathcal{A}(1)_{\mathbb{C}}^{\vee}}(B_{0}^{\mathbb{C}}(1))\implies\pi_{**}^{\mathbb{C}}(\text{ksp}).$
This Ext group was computed by the $\textbf{aAHSS}(B_{0}^{\mathbb{C}}(1))$ and is depicted in Fig. 5. There are no differentials for degree reasons, so the spectral sequence collapses and gives the result. ∎
Classically, one can express the groups $\text{Ext}_{\mathcal{A}(1)^{\vee}}(B_{0}(1)^{\otimes i})$ in terms of Adams covers of bo and bsp. This does not translate well to motivic homotopy theory. Over $\mathbb{C}$, Culver–Quigley [CQ21] show that one cannot express the groups $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(B_{0}^{\mathbb{C}}(1)^{\otimes i})$ in terms of Adams covers of kq and ksp. We show a similar result over $\mathbb{R}$ in Section 7. Instead, there is a periodic family of groups which will determine $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(B_{0}^{\mathbb{C}}(1))$ composed of Adams covers of bo, bsp, kq and ksp.
@@
    $Z^{\mathbb{C}}_{i}=\bigoplus_{j=0}^{i/2-1}\Sigma^{4j,2j}\mathbb{M}_{2}^{\mathbb{C}}[h_{0}]\oplus\Sigma^{2i,i}\text{Ext}_{\mathcal{A}(1)_{\mathbb{C}}^{\vee}}(\mathbb{M}_{2}^{\mathbb{C}})$
@@
    $Z^{\mathbb{C}}_{i}=\bigoplus_{j=0}^{(i-1)/2-1}\Sigma^{4j,2j}\mathbb{M}_{2}^{\mathbb{C}}[h_{0}]\oplus\Sigma^{2i-1,i-1}\text{Ext}_{\mathcal{A}(1)_{\mathbb{C}}^{\vee}}(B_{0}^{\mathbb{C}}(1));$
@@
    $Z^{\mathbb{C}}_{i}=\bigoplus_{j=0}^{i/2-1}\Sigma^{4j,2j}\mathbb{M}_{2}^{\mathbb{C}}[h_{0}]\oplus\Sigma^{2i-2,i}\mathbb{M}_{2}^{\mathbb{C}}\oplus\Sigma^{2i,i}\text{Ext}_{\mathcal{A}(1)_{\mathbb{C}}^{\vee}}(B_{0}^{\mathbb{C}}(1))\langle 1\rangle;$
@@
Proposition 5.5.
[CQ21, Lemma 3.36] There is an isomorphism of $\mathcal{A}(1)^{\vee}_{\mathbb{C}}$-comodules and $\textup{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(\mathbb{M}_{2}^{\mathbb{C}})$-modules:
$\frac{\textup{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(B_{0}^{\mathbb{C}}(1)^{\otimes i})}{v_{1}\textup{-torsion}}\cong Z_{i}^{\mathbb{C}}.$

### v2

@@ 4.2.2. Ext of higher integral Brown–Gitler comodules
$0\to\Sigma^{2(k-1),k-1}B_{0}^{F}(\tfrac{k-1}{2})\otimes_{\mathbb{M}_{2}^{F}}B_{0}^{F}(1)\to B_{0}^{F}(k)\to B_{1}^{F}(\tfrac{k-1}{2}-1)\otimes_{\mathbb{M}_{2}^{F}}(\mathcal{A}(1)\operatorname{/\!/}\mathcal{A}(0))_{F}^{\vee}\to 0.$
In the same way as before, the long exact sequence induced by $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{F}}(-)$ splits into short exact sequences modulo $v_{1}$-torsion. The cokernels are again a simple computation, but the kernels require a more delicate analysis. Indeed, to compute $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{F}}\left(B_{0}^{F}(\tfrac{k-1}{2})\otimes_{\mathbb{M}_{2}^{F}}B_{0}^{F}(1)\right)$, we use another aAHSS. This spectral sequence comes from applying $\text{Ext}^{***}_{\mathcal{A}(1)_{F}^{\vee}}\left(B_{0}^{F}(\tfrac{k-1}{2})\otimes_{\mathbb{M}_{2}^{F}}-\right)$ to (4.3). The $\textup{E}_{1}$-page is given by
$\textup{E}_{1}=\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}}(B_{0}^{F}(\tfrac{k-1}{2}))\otimes_{\mathbb{M}_{2}^{F}}\mathbb{M}_{2}^{F}\{[1],[\overline{\xi}_{1}],[\overline{\tau}_{1}]\}.$
@@
By 4.1, assembling these results gives the $\textup{E}_{2}$-page of the $\textbf{mASS}^{F}(\textup{kq}\otimes\textup{kq})$, modulo $v_{1}$-torsion. We can now hope to analyze this spectral sequence by comparing with the classical case and base change. In [17], this was shown to collapse on the $\textup{E}_{2}$-page for all algebraically closed fields of characteristic 0, such as $\mathbb{C}$. We will show that this is also true over $\mathbb{R}$.
@@ 5. Computations in $\text{SH}(\mathbb{C})$
In this section, we review Culver–Quigley’s [17] computation of $\pi_{**}^{\mathbb{C}}(\textup{kq}\otimes\textup{kq})$. This serves both as an example of the methods described in Section 2 and as a place to recall results which will be useful in the sequel via base change. There are a few places where our arguments differ from the ones given in [17]; we will explicitly indicate when this is the case.
@@ 5.1. Background
Let $\mathbb{M}_{2}^{\mathbb{C}}=\mathbb{F}_{2}[\tau]$ denote the mod-2 motivic homology of a point [55, Corollary 6.10], where $|\tau|=(0,-1).$ Recall that as $\rho=0$ in $\mathbb{M}_{2}^{\mathbb{C}}$, the dual Steenrod algebra is an honest Hopf algebra (see Remark 2.4). For the reader’s convenience, we review its structure.
Theorem 5.1 ([56, Theorem 12.6]).
The dual motivic Steenrod algebra $\mathcal{A}^{\vee}_{\mathbb{C}}=\pi_{**}^{\mathbb{C}}(\textup{H}\mathbb{F}_{2}\otimes\textup{H}\mathbb{F}_{2})$ is the commutative Hopf algebra given by
@@
There is a well understood relationship between $\mathcal{A}_{\mathbb{C}}^{\vee}$ and the classical dual Steenrod algebra $\mathcal{A}^{\vee}_{\text{cl}}$. Complex Betti realization gives a map $\text{Be}^{\mathbb{C}}:\mathcal{A}^{\vee}_{\mathbb{C}}\to\mathcal{A}^{\vee}_{\text{cl}}$ determined by
@@
This extends to an isomorphism in Ext [19]:
$\text{Ext}^{***}_{\mathcal{A}_{\mathbb{C}}^{\vee}}(\mathbb{M}_{2}^{\mathbb{C}})\otimes_{\mathbb{M}_{2}^{\mathbb{C}}}\mathbb{M}_{2}^{\mathbb{C}}[\tau^{-1}]\cong\text{Ext}^{**}_{\mathcal{A}_{\text{cl}}^{\vee}}(\mathbb{F}_{2})\otimes\mathbb{F}_{2}[\tau^{\pm 1}].$
This observation has proved very useful in $\mathbb{C}$-motivic computations [31, 30] and sits at the heart of the theory of synthetic spectra [22, 21, 48]. We will use this isomorphism to determine motivic information from classical information.
In addition to an understanding of the homology $\textup{H}_{**}(\textup{kq})\cong(\mathcal{A}\operatorname{/\!/}\mathcal{A}(1))_{\mathbb{C}}^{\vee}$, we also have a computation of the homotopy groups $\pi_{**}^{\mathbb{C}}(\textup{kq})$ [29].
Theorem 5.2 ([29, Theorem 4.10]).
The $\textup{{mASS}}^{\mathbb{C}}(\textup{kq})$ takes the form
$\textup{E}_{2}^{s,f,w}=\textup{Ext}^{s,f,w}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(\mathbb{M}_{2}^{\mathbb{C}})=\frac{\mathbb{M}_{2}^{\mathbb{C}}[h_{0},h_{1},a,b]}{(h_{0}h_{1},\tau h_{1}^{3},h_{1}a,a^{2}=h_{0}^{2}b)}\implies\pi_{s,w}^{\mathbb{C}}(\textup{kq}),$
where $|h_{0}|=(0,1,0)$, $|h_{1}|=(1,1,1)$, $|a|=(4,3,2)$ and $|b|=(8,4,4)$. The spectral sequence collapses at $\textup{E}_{2}$ with no extensions, giving an isomorphism
$\pi_{**}^{\mathbb{C}}(\textup{kq})\cong\frac{\mathbb{Z}_{2}^{\wedge}[\tau,\eta,\alpha,\beta]}{(2\eta,\tau\eta^{3},\eta\alpha,\alpha^{2}=4\beta)},$
@@
Following our discussion in Section 2, the first step towards computing the kq-resolution is computing the ring of cooperations $\pi_{**}^{\mathbb{C}}(\textup{kq}\otimes\textup{kq})$. Recall from 4.1 that the $\textbf{mASS}^{\mathbb{C}}(\textup{kq}\otimes\textup{kq})$ takes the form
$\textup{E}_{2}^{s,f,w}=\bigoplus_{k\geq 0}\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}^{s,f,w}(\Sigma^{4k,2k}B_{0}^{\mathbb{C}}(k))\implies\pi_{s,w}^{\mathbb{C}}(\text{kq}\otimes\text{kq}),$
where $B_{0}^{\mathbb{C}}(k)$ are the $\mathbb{C}$-motivic integral motivic Brown–Gitler comodules. Therefore, we must compute the trigraded groups $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(B_{0}^{\mathbb{C}}(k))$ for $k\geq 0.$ Since $B_{0}^{\mathbb{C}}(0)=\mathbb{M}_{2}^{\mathbb{C}}$, the first summand is computed by 5.2. We move then to compute $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(B_{0}^{\mathbb{C}}(1))$ by the algebraic Atiyah–Hirzebruch spectral sequence outlined in Section 4.2.
5.2. $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(B_{0}^{\mathbb{C}}(1)^{\otimes i})$
@@
$\textup{E}_{1}^{s,f,w,a}=\text{Ext}^{s,f,w}_{\mathcal{A}(1)_{\mathbb{C}}^{\vee}}(\mathbb{M}_{2}^{\mathbb{C}})\otimes_{\mathbb{M}_{2}^{\mathbb{C}}}\mathbb{M}^{\mathbb{C}}_{2}\{[1],[\overline{\xi}_{1}],[\overline{\tau}_{1}]\}\implies\text{Ext}^{s,f,w}_{\mathcal{A}(1)_{\mathbb{C}}^{\vee}}(B_{0}^{\mathbb{C}}(1)),$
with differentials
$d_{r}:\text{E}_{r}^{s,f,w,a}\to\text{E}_{r}^{s-1,f+1,w,a-r}.$
We will let $\alpha[i]$ denote the copy of a class $\alpha\in\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(\mathbb{M}_{2}^{\mathbb{C}})$ in Atiyah–Hirzebruch filtration $i$. We have discussed in Section 4.2 the differentials for this spectral sequence. Namely, the $d_{1}$-differential is only nonzero on Atiyah–Hirzebruch filtration 3, where
@@
There is a potential nonzero third differential $d_{3}:\textup{E}_{3}^{s,f,w,3}\to\textup{E}_{3}^{s-1,f+1,w,0}$. However, for degree reasons we have that $\langle\alpha,h_{0},h_{1}\rangle=0$ for all remaining classes $\alpha[3]\in\textup{E}_{3}^{s,f,w,3}$, hence this differential is 0. Thus the spectral sequence collapses on $\textup{E}_{3}=\textup{E}_{\infty}$.
We depict the $\textup{E}_{1}$-page in Figure 5.1, the $\textup{E}_{2}$-page in Figure 5.2, and the $\textup{E}_{3}=\textup{E}_{\infty}$-page in Figure 5.3. Charts are written in $(s,f)$-grading, with motivic weight and Atiyah–Hirzebruch filtration suppressed. A black $\bullet$ represents $\mathbb{M}_{2}^{\mathbb{C}}$. A hollow $\circ$ represents $\mathbb{F}_{2}$. A black vertical line represents multiplication by $h_{0}$. A diagonal line represents multiplication by $h_{1}$. Differentials are blue, linear with respect to the $\text{Ext}^{***}_{\mathcal{A}(1)_{\mathbb{C}}^{\vee}}(\mathbb{M}_{2}^{\mathbb{C}})$-module structure, and preserve motivic weight. A dashed blue line represents a differential where either source or target is $\tau$-torsion.
Figure 5.1. The $\textup{E}_{1}$-page of the $\textbf{aAHSS}(B_{0}^{\mathbb{C}}(1))$
@@
[…]
The hidden extensions on the $\textup{E}_{\infty}$-page can be resolved using complex Betti realization. Recall from Section 2.1 that this is a functor $\text{Be}^{\mathbb{C}}:\text{SH}(\mathbb{C})\to\text{Sp}$. For $X\in\text{SH}(\mathbb{C})$, this gives a map in homotopy
$\pi_{s,w}^{\mathbb{C}}(X)\to\pi_{s}(\text{Be}^{\mathbb{C}}(X)).$
In particular, there is a map of dual Steenrod algebras $\mathcal{A}^{\vee}_{\mathbb{C}}\to\mathcal{A}_{\text{cl}}^{\vee}$ sending $\bar{\xi}_{i}\mapsto\bar{\xi}_{i}^{2}$ and $\bar{\tau}_{i}\mapsto\bar{\xi}_{i}$. Additionally, there is an isomorphism of $\mathcal{A}(1)_{\mathbb{C}}^{\vee}$-comodules
$(B_{0}^{\mathbb{C}}(1))[\tau^{-1}]\cong B_{0}^{\text{cl}}(1)[\tau^{\pm 1}],$
where $B_{0}^{\text{cl}}(1)$ is the classical Brown–Gitler comodule. This induces a map of Ext groups
$\text{Ext}^{s,f,w}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(B_{0}^{\mathbb{C}}(k))\to\text{Ext}_{\mathcal{A}(1)_{\text{cl}}^{\vee}}^{s,f}(B_{0}(k)),$
and allows one to lift hidden extensions from the classical case in [41] to the $\mathbb{C}$-motivic case.
@@
Corollary 5.3 ([17, Corollary 3.31]).
The $\textup{{mASS}}^{\mathbb{C}}(\textup{ksp})$ has signature
$\textup{E}_{2}^{s,f,w}=\textup{Ext}^{s,f,w}_{\mathcal{A}(1)_{\mathbb{C}}^{\vee}}(B_{0}^{\mathbb{C}}(1))\implies\pi^{\mathbb{C}}_{s,w}(\textup{ksp})$
and collapses on the $\textup{E}_{2}$-page.
@@
A change of rings isomorphism gives the $\textup{E}_{2}$-page for the $\textbf{mASS}^{\mathbb{C}}(\text{ksp})$. This Ext group was computed by the $\textbf{aAHSS}(B_{0}^{\mathbb{C}}(1))$ and is depicted in Figure 5.3. There are no differentials for degree reasons, so the spectral sequence collapses and gives the result. ∎
Classically, one can express the groups $\text{Ext}^{**}_{\mathcal{A}(1)_{\text{cl}}^{\vee}}(B_{0}(1)^{\otimes i})$ in terms of Adams covers of bo and bsp. This does not translate well to motivic homotopy theory. Over $\mathbb{C}$, Culver–Quigley [17] show that one cannot express the groups $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(B_{0}^{\mathbb{C}}(1)^{\otimes i})$ in terms of Adams covers of kq and ksp. We show a similar result over $\mathbb{R}$ in Section 7. Instead, there is a periodic family of groups which will determine $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(B_{0}^{\mathbb{C}}(1))$ composed of Adams covers of bo, bsp, kq and ksp.
@@
    $Z^{\mathbb{C}}_{i}=\bigoplus_{j=0}^{i/2-1}\Sigma^{4j,2j}\mathbb{M}_{2}^{\mathbb{C}}[h_{0}]\oplus\Sigma^{2i,i}\text{Ext}^{***}_{\mathcal{A}(1)_{\mathbb{C}}^{\vee}}(\mathbb{M}_{2}^{\mathbb{C}})$
@@
    $Z^{\mathbb{C}}_{i}=\bigoplus_{j=0}^{(i-1)/2-1}\Sigma^{4j,2j}\mathbb{M}_{2}^{\mathbb{C}}[h_{0}]\oplus\Sigma^{2i-1,i-1}\text{Ext}^{***}_{\mathcal{A}(1)_{\mathbb{C}}^{\vee}}(B_{0}^{\mathbb{C}}(1));$
@@
    $Z^{\mathbb{C}}_{i}=\bigoplus_{j=0}^{i/2-1}\Sigma^{4j,2j}\mathbb{M}_{2}^{\mathbb{C}}[h_{0}]\oplus\Sigma^{2i-2,i}\mathbb{M}_{2}^{\mathbb{C}}\oplus\Sigma^{2i,i}\text{Ext}^{***}_{\mathcal{A}(1)_{\mathbb{C}}^{\vee}}(B_{0}^{\mathbb{C}}(1))\langle 1\rangle;$
@@
Proposition 5.4 ([17, Lemma 3.36]).
There is an isomorphism of $\mathcal{A}(1)^{\vee}_{\mathbb{C}}$-comodules and $\textup{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(\mathbb{M}_{2}^{\mathbb{C}})$-modules:
$\frac{\textup{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(B_{0}^{\mathbb{C}}(1)^{\otimes i})}{v_{1}\textup{-torsion}}\cong Z_{i}^{\mathbb{C}}.$